OpenStudy (anonymous):

show that the function x^3-6x^2+12x+9 is increasing on all interval

OpenStudy (matt101):

To find out whether the function is increasing or decreasing, first you need to find the critical points. What are the critical points of this function?

OpenStudy (anonymous):

i know that but the question is , to show that it is increasing at all intervals

OpenStudy (anonymous):

how do i go about that?

OpenStudy (matt101):

Well after you have the critical points, you'll have a bunch of intervals from one critical point to the next. All you need to do is plug in any x-value within each interval into f'(x) and see what you get (because all the x-values within a given interval will follow the same trend in terms of slope). If f'(x) > 0 (the slope is positive), then the interval is increasing. If f'(x) < 0 (the slope is negative), then the interval is decreasing. If f'(x) > 0 for every interval, then the function is always increasing.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Since you have to show that the function is increasing everywhere, you only need to establish that its derivative is positive. \[f(x)=x^3-6x^2+12x+9~~\implies~~f'(x)=3x^2-12x+12=3(x-2)^2\] What do you know about anything that gets squared?

OpenStudy (anonymous):

waw. thats cool. all the values are positive

OpenStudy (anonymous):

Well, everywhere except when \(x=6\).

OpenStudy (perl):

@SithsAndGiggles everywhere except x = 2 (just checking)

OpenStudy (anonymous):

Right, I was thinking of a different question. Thanks